OpenStudy (anonymous):

Please help x-4/5=6/13

rebeccaxhawaii (rebeccaxhawaii):

add -4/5 to both sides

OpenStudy (anonymous):

you should add +4/5 to the both sides

OpenStudy (jdoe0001):

\(\bf x-\cfrac{4}{5}=\cfrac{6}{13}\qquad or\qquad \cfrac{x-4}{5}=\cfrac{6}{13}?\)

OpenStudy (anonymous):

Oh, 2nd one

OpenStudy (jdoe0001):

then just cross-multiply and siimplify

OpenStudy (anonymous):

Ok, can you help me with that?

OpenStudy (jdoe0001):

\(\bf \cfrac{x-4}{{\color{blue}{ 5}}}=\cfrac{6}{{\color{brown}{ 13}}}\implies {\color{brown}{ 13}}(x-4)={\color{blue}{ 5}}(6)\) expand both sides

OpenStudy (anonymous):

Ok, what do I do about (x-4) though?

OpenStudy (jdoe0001):

distribute the 13 to it :/

OpenStudy (anonymous):

Alright, thanks.

OpenStudy (jdoe0001):

yw